OpenStudy (anonymous):

For what value of a, would the system have an infinite number of solutions? 2x-y=5 8x-47=2a A) 2 B) 4 C)5 D)10 E)20

OpenStudy (tkhunny):

Pick the value that makes them identical in every way. You have: 2x-7=5 8x-47=2a A little algebra 2x = 5 + 7 = 12 8x = 2a + 47 Multiply the top one by 4. 8x = 48 8x = 2a + 47 Hmmm... Now what?

OpenStudy (anonymous):

@tkhunny i wrote the problem wrong 7 is supposed to be y

OpenStudy (tkhunny):

Okay, use the same methodology to solve the problem.